The sum of two numbers is 23. Twice the greater number added to three times the smaller number is 57.
Answer by guapa(62)

Let x be the smaller number then y is the greater.
The sum (+) of two numbers (x, y) is (=) 23. Thus, x+y=23 This is our first equation.
You can solve this equation for x or y. Let's do it for x. x=23-y
Now, twice the greater number (2y) added (+) to three times the smaller number (3x)is 57 (=57)
So, the second equation is: 2y + 3x = 57
Now substitute (23-y) for x:
2y + 3(23-y)=57
2y+ 69-3y= 57
69-y=57
y=12
If y =12 then x=23-y equals x=23-12, x= 11
Now we can check our answers: x+y = 23, 11+12=23
And 2y+3x = 57, 2(12)+3(11)=57, 24+33=57